OpenStudy (anonymous):

Which expression is the simplified form of the square root 63 times w to the thirty-sixth power.? 3w6the square root of 7. 9w6the square root of 7. 3w18the square root of 7. 9w18the square root of 7.

OpenStudy (cwrw238):

sqrt (63 * w^36) sqrt63 = sqrt7*sqrt9 = 3 sqrt7 sqrt w^36 = w^18